OpenStudy (anonymous):

how does the product rule apply for the derivative of two variables say: z(x,y)= xy

OpenStudy (kmeis002):

Is this a class where you know partial derivatives?

OpenStudy (anonymous):

i'm not sure if i have to apply partial derivatives, i have to find the derivative of r(p) = pd(p) to get the elasticity

OpenStudy (kmeis002):

Okay, so does the question hint at assuming that y is a function of x? \[y(x) \]

OpenStudy (anonymous):

they use d(p) and r(p) and p so i guess they do

myininaya (myininaya):

well if you said use product rule for r(p)=p*d(p) then I would not think to use partial derivative but your other question does require parital derivatives

OpenStudy (anonymous):

question says: if d(p) represents the demand function, then the revenue r is equal to the product of the price p and the demand d(p), r(p)= pd(p). the elasticity of the revenue is denoted by er and the elasticity of the demand by ed. show that the following relation holds er= 1+ ed

OpenStudy (kmeis002):

Ah, okay, I though that was for a different question. \[ r(p) = pd(p)\] Apply the dervative operator w.r.t. p \[\frac{d}{dp} \left ( r(p) \right ) =\frac{d}{dp} \left ( p d(p) \right ) \] From the product rule, we can simplify the rhs to: \[\frac{d}{dp} \left ( r(p) \right ) =\frac{d}{dp}(p) d+ \frac{d}{dp}(d) p \] Since p and d are functions of p, we get: \[\frac{d}{dp} \left ( r(p) \right ) =d+ \frac{dd}{dp}p \]